what are prime numbers ;54,12,36
Answer by checkley71(8403)

prime numbers are numbers that can only be factored into 1 & the number it self.
EX: 5=1*5, 13=1*13, 37=1*37, etc.
these numbers are not prome:
54=1*2*27
12=1*2*2*3
36=1*2*2*3*3